# Fuzzy Prediction Sets: Conformal Prediction with E-values

### v2

@@ A.2 Proof of Proposition 1
If $Z^{n+1}$ is exchangeable, then $Z^{n}$ is exchangeable. Exchangeability of $Z^{n}$ means that the distribution of $Z^{n+1}$ is invariant under permutations in $\Pi(n)$, which is equivalent to uniformity on each $\Pi(n)$-orbit of $\mathcal{Z}^{n+1}$, as $\Pi(n)$ is a compact group. Each $\Pi(n)$-orbit of $\mathcal{Z}^{n+1}$ is simply the subset of a $\Pi(n+1)$-orbit $O\in\mathcal{O}^{n+1}$ in which the $(n+1)$th observation is fixed. In fact, each $\Pi(n+1)$-orbit may be partitioned into $\Pi(n)$-orbits as subsets, each differing by the $(n+1)$th element. Hence, conditionally on $O\in\mathcal{O}^{n+1}$, each conditional distribution $Q^{|O}$ may be viewed as first sampling this $(n+1)$th element using some distributions (and with it the $\Pi(n)$-orbit), and subsequently sampling uniformly from the $\Pi(n)$-orbit that contains this element. ∎
@@ A.3 Proof of Proposition 2
By Proposition 3 in koning2023post, the conditional likelihood ratio is the restriction of
@@
to $O$. By definition of $\overline{Q}$, we have, for every event $A$,

### v3

@@ A.3 Proof of Proposition 1
By assumption, $Z^{n}$ is exchangeable. Exchangeability of $Z^{n}$ means that the distribution of $Z^{n+1}$ is invariant under permutations in $\Pi(n)$, which is equivalent to uniformity on each $\Pi(n)$-orbit of $\mathcal{Z}^{n+1}$, as $\Pi(n)$ is a compact group. Each $\Pi(n)$-orbit of $\mathcal{Z}^{n+1}$ is simply the subset of a $\Pi(n+1)$-orbit $O\in\mathcal{O}^{n+1}$ in which the $(n+1)$th observation is fixed. In fact, each $\Pi(n+1)$-orbit may be partitioned into $\Pi(n)$-orbits as subsets, each differing by the $(n+1)$th element. Hence, conditionally on $O\in\mathcal{O}^{n+1}$, each conditional distribution $Q^{|O}$ may be viewed as first sampling this $(n+1)$th element using some distributions (and with it the $\Pi(n)$-orbit), and subsequently sampling uniformly from the $\Pi(n)$-orbit that contains this element. ∎
@@ A.4 Proof of Proposition 2
By Proposition 3 in Koning (2025b), the conditional likelihood ratio is the restriction of
@@
to $O$. By definition of $\overline{Q}$, we have, for every event AA,
@@
A.5 Proof of Proposition 3
Proof.
We have the implication $L_{Z_{n+1}}(\delta)>\sup_{z\in C_{\alpha}^{Z^{n}}}L_{z}(\delta)\implies Z_{n+1}\not\in C_{\alpha}^{Z^{n}}$, so that $P\left(L_{Z_{n+1}}(\delta)>\sup_{z\in C_{\alpha}^{Z^{n}}}L_{z}(\delta)\right)\leq P\left(Z_{n+1}\not\in C_{\alpha}^{Z^{n}}\right)\leq\alpha$. ∎
A.6 Proof of Proposition 5
Proof.
We have $\sup_{z\in\mathcal{Z}}L_{z}(\delta)/\mathcal{E}^{Z^{n}}(z)\geq L_{Z_{n+1}}(\delta)/\mathcal{E}^{Z^{n}}(Z_{n+1})$, so that
$\frac{L_{Z_{n+1}}(\delta)}{\sup_{z\in\mathcal{Z}}L_{z}(\delta)/\mathcal{E}^{Z^{n}}(z)}\leq\frac{L_{Z_{n+1}}(\delta)}{L_{Z_{n+1}}(\delta)/\mathcal{E}^{Z^{n}}(Z_{n+1})}=\mathcal{E}^{Z^{n}}(Z_{n+1}).$
The result then follows from the validity of $\mathcal{E}^{Z^{n}}$. ∎
Figure 6: Optimal fuzzy prediction sets for Neyman–Pearson-utility (solid), log-utility (dotted) and bounded log-utility (dashed) under the simple Gaussian setting (left) and composite Gaussian setting (right) from Example 1, 2 and 8, for $n=3$, $\mu=0$, $\sigma=1$, $\tau=3.5$, $\bar{Z}_{n}=1.44$. Here, the Neyman–Pearson utility and bounded log-utility are both bounded at $1/0.05=20$.
Appendix B Example Fuzzy Gaussian
Example 8 (Fuzzy Gaussian).
In the examples of Section 2.6, we implicitly considered optimal fuzzy prediction sets under the classical Neyman–Pearson utility $x\mapsto x\wedge 1/\alpha$. We continue these examples under more general utility functions.
We first consider the log-utility version of Example 1, where we had the simple model $\mathcal{P}=\{\mathcal{N}(\mu 1_{n+1},\sigma^{2}I_{n+1})\}$ and optimized against the alternative $Q=\mathcal{N}(\mu 1_{n},\sigma^{2}I_{n})\times\mathcal{N}(\mu,\tau^{2})$, for known $\mu\in\mathbb{R}$ and $\tau>\sigma>0$. Here, we could choose $\tau\to\infty$ to mimic the Lebesgue measure, as in the preceding examples, but we find that this generally yields undesirable fuzzy prediction sets.
For $1/\alpha$-bounded log-utility $U:x\mapsto\log(\mathcal{E}\wedge 1/\alpha)$, the optimal e-value is the likelihood ratio capped at $1/\alpha$ and ‘boosted’ by some constant $b_{\alpha}$ so that it has expectation exactly 1 (Koning, 2025a):
$\textnormal{LR}(A,B)=\frac{d\mathcal{N}(0,\tau^{2}+\sigma^{2}/n)}{d\mathcal{N}(0,\sigma^{2}+\sigma^{2}/n)}(B),$
$\mathcal{E}^{Z^{n}}(z_{n+1})=\frac{d\mathcal{N}(\bar{Z}_{n},\tau^{2}+\sigma^{2}/n)}{d\mathcal{N}(\bar{Z}_{n},\sigma^{2}+\sigma^{2}/n)}(z_{n+1}),$ (15)
which relies on a variant of the Hunt-Stein theorem derived for log-optimal e-values proven by Pérez-Ortiz et al. (2024). To the best of our knowledge, this has not been proven yet for general expected-utility-optimal e-values, but we would be surprised if it does not go through in general.
Comparing the fuzzy prediction set in (15) to (14), we see that this likelihood ratio is centered at $\bar{Z}_{n}$ instead of $\mu$ and the variance in both numerator and denominator is inflated by an additional $\sigma^{2}/n$-term due to the estimation of $\mu$ by $\bar{Z}_{n}$. We illustrate this fuzzy prediction set in the right panel of Figure 6, along its bounded version and the Neyman–Pearson-utility variant from Example 2.
Appendix C Clipping, capping and dampening
Analogously, in case we are not interested in evidence beyond some value $c\geq 0$, we may also cap our e-values below some upper bound $c\geq 1$. The resulting optimizer is analogous to (16), with $\vee b$ is replaced by $\wedge c$ (Koning, 2025a). Capping may also be incorporated in the utility function, and appears in the Neyman–Pearson utility function $U(x)=x\wedge 1/\alpha$.
An alternative strategy to ‘dampen’ e-values is introduced by Grünwald (2023). He proposes to take an e-value $\mathcal{E}$ and dampen it to the e-value $b+(1-b)\mathcal{E}$ that may be interpreted as a $b$-mixture of the e-value and the constant 1. If $\mathcal{E}$ is the log-utility maximizer, then we find $b+(1-b)\mathcal{E}$ implicitly maximizes the utility function $U(\mathcal{E})=\log([\mathcal{E}-b]\vee 0)$.
Appendix D Connecting as-if and weighted decisions
While the weighted loss approach for fuzzy prediction sets in Section 5.2.2 and the as-if decision approach from Section 5.1 may seem distinct, we show that the as-if approach may be viewed as a special limiting case of the weighted loss approach. This relationship was not yet established before; Andrews and Chen (2025) describe both approaches as distinct paradigms. The connection is not obvious, as it relies on our observation that the E-posterior coined by Grünwald (2023) is equivalent to a fuzzy confidence set, and so a generalization of a confidence set.
which is proportional to the risk function we minimize for non-fuzzy prediction sets in (9). This shows that the non-fuzzy ‘as-if’ decision can indeed be viewed as a limiting case of the fuzzy procedure.
Appendix E Covariates
$P(Y_{n+1}\in C_{\alpha}^{(Z^{n},X_{n+1})})\geq 1-\alpha,$
for every $P\in\mathcal{P}$.
$Q=P_{*}^{Z^{n},X_{n+1}}\otimes\mu^{|(Z^{n},X_{n+1})}.$
E.1 Conformal prediction with covariates
For example, the i.i.d. result in Proposition 2 now specializes to
$\textnormal{LR}^{|O(z^{n+1})}(z^{n+1})$ $\propto\frac{dQ^{Z_{n+1}}}{dQ^{Z_{1}}}(z^{n+1})=\frac{d(P_{*}^{X_{n+1}}\otimes\mu^{|X_{n+1}})}{dP_{*}^{X_{n+1},Y_{n+1}}}(z_{n+1})=\frac{dP_{*}^{X_{n+1}}\otimes d\mu^{|X_{n+1}}}{dP_{*}^{X_{n+1}}\otimes dP_{*}^{Y_{n+1}|X_{n+1}}}(z_{n+1})$
$=\frac{d\mu^{|X_{n+1}}}{dP_{*}^{Y_{n+1}|X_{n+1}}}(y_{n+1}),$
where the normalization constant is
$\frac{1}{(n+1)}\sum_{i=1}^{n+1}\frac{d\mu^{|X_{i}}}{dP_{*}^{Y_{n+1}|X_{i}}}(y_{i}).$